Churchill Products is considering updating its cost system to an activity-based costing system and is interested in understanding the effects. The company’s cost accountant has identified three overhead cost pools along with appropriate cost drivers for each pool.

Cost Pools Costs Activity Drivers
Utilities $ 320,000 64,000 machine-hours
Scheduling and setup 310,000 620 setups
Material handling 795,000 1,590,000 pounds of material

The company manufactures three models of water basins (Oval, Round, and Square). The plans for production for the next year and the budgeted direct costs and activity by product line are as follows.

Products
Oval Round Square
Total direct costs (material and labor) $ 80,000 $ 80,000 $ 90,000
Total machine-hours 30,000 12,000 22,000
Total number of setups 80 320 220
Total pounds of material 540,000 270,000 780,000
Total direct labor-hours 3,000 2,000 5,000
Number of units produced 4,000 2,000 6,000

Required:

a. The current cost accounting system charges overhead to products based on direct labor-hours. What unit product costs will be reported for the three products if the current cost system continues to be used?
c. What are the cost driver rates for the three cost pools identified by the cost accountant?
d. What unit product costs will be reported for the three products if the ABC system suggested by the cost accountant's classification of cost pools is used?
e. If management should decide to implement an activity-based costing system, what benefits should it expect?
Answer #1

Answer

a Total direct labour hours

= 3,000 + 2,000 + 5,000 = 10,000 hours

Total overhead

= 320,000 + 310,000 + 795,000 = $1,425,000

Overhead rate under current costing system

= 1,425,000 / 10,000 = $142.5

OVERHEAD ASSIGNED TO EACH PRODUCTS

To oval = 3,000 × 142.5 = $427,500

To Round = 2,000 × 142.5 = $285,000

To Square = 5,000 × 142.5 = $712,500

TOTAL COST OF EACH PRODUCTS

Oval = 80,000 + 427,500 = $507,500

Round = 80,000 + 285,000 = $365,000

Square = 90,000 + 712,500 = $802,500

UNIT PRODUCT COST OF EACH PRODUCTS UNDER CURRENT COSTING SYSTEM

Product = total cost / number of units produced

Oval = 507,500 / 4,000 = $126.875

Round = 365,000 / 2,000 = $182.5

Square = 802,500 / 6,000 = $133.75

b COST DRIVER RATES OF EACH COST POOLS

Utilities = 320,000 / 64,000 = $5

Schedule and setup = 310,000 / 620 = $500

Material handling = 795,000 / 1,590,000 = $0.5

c OVERHEAD ASSIGNED TO EACH PRODUCTS

Oval

= (30,000 × 5) + (80 × 500) + (540,000 × 0.5) = $460,000

Round

= (12,000 × 5) + (320 × 500) + (270,000 × 0.5) = $355,000

Square

= (22,000 × 5) + (220 × 500) + (780,000 × 0.5) = $610,000

TOTAL COST OF EACH PRODUCTS

Oval = 80,000 + 460,000 = $540,000

Round = 80,000 + 355,000 = $435,000

Square = 90,000 + 610,000 = $700,000

UNIT COST OF EACH PRODUCT UNDER ABC COSTING

Oval = 540,000 / 4,000 = $135

Round = 435,000 / 2,000 = $217.5

Square = 700,000 / 6,000 = $116.67

d The ABC costing system give us the more accurate allocation of overhead cost to product based on each product consumption of overhead resources. We get more accurate product cost compared to other costing systems.

For example oval unit product cost under current system is $126.875 and under ABC , it is $135.

$135 is the more unit product cost.

If YOU are like the answer please upvote.